OpenStudy (anonymous):

sqrt(x-3)-sqrt(2x+1)=3

OpenStudy (adunb8):

are you looking for x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

square both sides [swrt(x-3)]-[sqrt(2x+1)] = 3^2 Get x-3-2x+1 = 9 (x-2x) + (-3+1) = 9 -x -2 = 9 Add 2 -x= 11 To get x by itself, and positive, divide by -1 x=-11

OpenStudy (anonymous):

I tried replacing x with that value, it doesnt work

OpenStudy (dumbcow):

notice a couple errors...first of all always check your answers, if you plug in x=-11 you get sqrt of neg number thus it is not a solution also, you didn't square left side properly...you have to FOIL

OpenStudy (anonymous):

ORLY? okay, let me try again. I like that username, 'dumbcow' kudos to you.

OpenStudy (anonymous):

OMGUSH. foil. FOIL. Shicuss me out of here, i suck

OpenStudy (dumbcow):

haha..yep the best username

OpenStudy (anonymous):

okay, okay, so then help us out. what exactly are you foiling. the x-3 and 2x + 1?

OpenStudy (dumbcow):

\[[\sqrt{x-3} -\sqrt{2x+1}]^{2} = (x-3) -2 \sqrt{x-3}\sqrt{2x+1}+(2x+1)\]

OpenStudy (anonymous):

AH. the same mistake i always make. Y'know, I made that same mistake on my math test two years ago.

OpenStudy (dumbcow):

common mistake with students... \[(a-b)^{2} \ne a^{2} -b^{2}\]

OpenStudy (anonymous):

dude, how do you get that fancy stuff? :I you're making me jelly. and hold on, I shall do this over again. FAILURE IS NOT AN OPTION.

OpenStudy (dumbcow):

@erick193 , are you still following? this is kinda a long problem with lot of steps

OpenStudy (dumbcow):

once you square both sides...next step is to isolate radical(s) on 1 side

OpenStudy (anonymous):

Ahahhahaha. okay, jk, momma yelling at me right now, brb. if im not back, i hope you find the answer erick

OpenStudy (anonymous):

and how you do that?

OpenStudy (dumbcow):

\[(x-3)-2\sqrt{x-3} \sqrt{2x+1}+(2x+1) = 9\] (x-3) + (2x+1) = 3x-2 right? \[-2\sqrt{x-3}\sqrt{2x+1} +3x-2 = 9\] then move the 3x-2 to other side \[-2\sqrt{x-3}\sqrt{2x+1} = 11-3x\] divide by -2 \[\sqrt{x-3}\sqrt{2x+1} = \frac{11-3x}{-2}\]

OpenStudy (anonymous):

so how do i get rid of radicals?

OpenStudy (dumbcow):

next step is square both sides again....always square both sides to get rid of radicals this doesn't come out nice, are you sure the question is written correctly

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I just asked a classmate on FB and he says there is no solution

OpenStudy (anonymous):

I also tried it and only keep getting radicals after I square

OpenStudy (dumbcow):

yes, there is no Solution....both answers are extraneous but you still need to understand the steps so you can do other similar problems when you square a radical , it goes away by definition \[(\sqrt{x-3}\sqrt{2x+1})^{2} = (x-3)(2x+1)\]

OpenStudy (anonymous):

ok..

OpenStudy (dumbcow):

the faster, easier way to check for No solution is graph it with calculator http://www.wolframalpha.com/input/?i=sqrt%28x-3%29-sqrt%282x%2B1%29+%3D+3 notice the curve from left side never intersects with line y=3

OpenStudy (anonymous):

So if the lines never intersect, it means no solution?

OpenStudy (dumbcow):

correct...solutions are points where 2 lines intersect

OpenStudy (anonymous):

Is there a way to know if there is no solution without a graph calculator?

OpenStudy (dumbcow):

hmm well you could graph it by hand but like for this problem that would be too hard. otherwise...you have to go through and solve for x, then plug them into original equation to see if they work, if None work then No solution

OpenStudy (anonymous):

Ok thankyou!

OpenStudy (dumbcow):

your welcome